In Drosophila, the genes for black (b) body and curved (c) wings are linked on the same chromosome. The frequency of recombination between b---c is 0.276. How many recombinant offspring would be expected if 2000 progeny were produced from a testcross between a female heterozygous for both genes and a homozygous recessive male?
Answer #1

As it is said that test cross produced 2000 progeny and the recombination frequency is 0.276. So, according to the formula,

Recombination frequency= recombinant progenies from test cross/ total no. Of progenies.

So here recombination frequency is given and total no. Of progenies is also given, that mean no. Of recombinant offspring can be calculated,

0.276 = x / 2000

X = 0.276*2000 = 552

So there will be total 552 recombinant progenies.